# Sakib Matina, Alice Allena,b, Emily Shinklec, Aleksandra Pachalievad, Galen T. Cravena, Benjamin Nebgena, Justin Smithe, Richard Messerlya, Ying Wai Lib, Sergei Tretiaka,b,f, Kipton Barrosa,b, and Nicholas Lubbersc

### v1

@@
[Uncaptioned image] Teacher-student training improves accuracy and efficiency of machine learning inter-atomic potentials
Sakib Matina, Alice Allena,b, Emily Shinklec, Aleksandra Pachalievad, Galen T. Cravena, Benjamin Nebgena, Justin Smithe, Richard Messerlya, Ying Wai Lib, Sergei Tretiaka,b,f, Kipton Barrosa,b, and Nicholas Lubbersc
[Uncaptioned image] Machine learning inter-atomic potentials (MLIPs) are revolutionizing the field of molecular dynamics (MD) simulations. Recent MLIPs have tended towards more complex architectures trained on larger datasets. The resulting increase in computational and memory costs may prohibit the application of these MLIPs to perform large-scale MD simulations. Here, we present a teacher-student training framework in which the latent knowledge from the teacher (atomic energies) is used to augment the students’ training. We show that the light-weight student MLIPs have faster MD speeds at a fraction of the memory footprint compared to the teacher models. Remarkably, the student models can even surpass the accuracy of the teachers, even though both are trained on the same quantum chemistry dataset. Our work highlights a practical method for MLIPs to reduce the resources required for large-scale MD simulations.
@@ 1 Introduction
Molecular Dynamics (MD) 1 is ubiquitous in chemistry 2, materials science 3, and drug discovery 4 as well as other fields. Accurate chemical and thermodynamic properties derived from MD rely on accurate inter-atomic potentials, which parameterize the many-body interactions present between atoms 1, 5. Simulation scales may vary greatly depending on the questions of interest and available resources, and there is a persistent need for greater model efficiency. Traditional classical potentials are very fast, making it possible to perform large-scale simulations of billions of atom 6, or to perform hundreds of millions of integration time-steps for small systems. Recently, there is great demand for interatomic potentials that are more accurate via machine learning models that are trained to reference quantum mechanical force calculations. Here, the goals of efficiency and accuracy can be in conflict. Our paper is concerned with techniques for improving the efficiency of the inter-atomic potential without sacrificing accuracy.
@@
The gold-standard for computational chemistry is ab initio molecular dynamics, which uses quantum chemistry (QC) methods for accurately calculating inter-atomic forces from first principles 7, 8. However, the often prohibitive computational cost of QC methods, and its rapid growth with system size, limits the size of the systems that can be simulated. Machine learning inter-atomic potentials (MLIPs) 9, 10, 11, 12, 13 can be trained on QC datasets to map from an atomic configuration to energy and forces. MLIPs can achieve the chemical accuracy (error $<1\mathrm{kcal/mol}$) of QC simulations 14 at drastically reduced computational costs. Most MLIPs achieve linear scaling with system size by utilizing the approximation that the total predicted energy can be decomposed as a sum of spatially local atom-wise or pair-wise contributions 15, 12, 11, 16. MLIPs have seen explosive growth and have been successfully applied to predicting potential energy surfaces 17, 18, 19, 20, 21, 22, 23, 24, 25, with extensions to a variety of quantities, such as charges 26, 27, 28, 29, 30, and more 31, 32, 33, 34, 35. However, MLIPs only mitigate the cost of QC—they do not eliminate it. Generating training data for MLIPs is costly due to the steep scaling of QC methods in system size $N$, e.g., $\mathcal{O}(N^{7})$ at the CCSD(T) level of theory 7. Therefore any approach which uses data more efficiently can potentially reduce the overall computational costs of MLIPs.
@@
Furthermore, a recent trend in the field is the significant effort to construct foundation model MLIPs 36, 37, 38 with broad chemical coverage39, inspired by the success of large pre-trained models in natural language processing 40. Foundation MLIPs 41, which may be parameterized by up to $10^{9}$ fitting parameters 42, have high computational and memory requirements 43. These increasing computational and memory costs compared to classical force fields 1 can limit the usability of MLIPs for large scale MD simulations 13, 44, 43.
@@
In this manuscript, we introduce a teacher-student training method for MLIP with a central purpose of building the MLIPs with faster inference and lower memory requirements. The teacher-student training is a class of methods 45, 46, 47 to reduce the inference costs of different ML models and more effectively utilize existing datasets. An initial teacher model is trained and then used to augment the training of a student model that may have faster inference 47, 43, smaller memory requirements 48, or better generalization capacity 46. Crucially, this knowledge distillation procedure does not require additional first-principles training data. Instead, auxiliary predictions of the teacher model are used to augment the data used for training of the student model. The innovation in this work is to use the teacher’s local atomic energy predictions as the auxiliary training data. Although these local atomic energies have traditionally been considered a latent feature of the MLIP more 11, the present work highlights that they carry important information. Note that the latent atomic energies provided to the student are far greater in number than the single global QC energy. Thus the student is trained using a significantly larger number of constraints than the teacher. Because the student model will typically have fewer trainable weighs than the teacher model, this approach can yield significant gains for inference speed and memory requirements. We also find that the student model can achieve higher-accuracy than the teacher. An overview of our workflow is represented in Fig 1.
@@
Fig.  1: Teacher-student training for Hierarchically Interacting Particle Neural Network (HIPNN) machine learning inter-atomic potential. A HIPNN model consists of an input node, a message passing interaction layer and feed-forward regression layers called atom layers. The teacher HIPNN is trained on the quantum mechanical energy and forces data, and generates latent knowledge in the form of atomic energies. This augments the student model training to improve the accuracy. The student HIPNN that has fewer trainable weights, which contributes to faster inference and lower memory requirements. The control HIPNNs have the same number of trainable parameters as the student but are trained only on ground truth data. We show that the student models are more accurate than the control models.
@@ 2.1 Architecture
HIPNN 22 is a message-passing graph neural network 49 that can map atomic configurations to various chemical properties such as energy 22, forces 14, dipoles 28 etc. Similar to many MLIP architectures, the interaction layers of HIPNN allow for mixing of atomic environments between neighbors via message passing to construct the learnable features 22. The local atomic environment of each atom is initially featurized using atomic number (in HIPNN, as a one-hot encoding, although also common is a random embedding) and passed through several layers along with pair-wise displacement vectors between the neighbors with local cut-off to predict the atom energy $\mathcal{\epsilon}_{i}$ for each atom $i$. Automatic differentiation is used to calculate the forces on each atom from the total molecular energy.
@@
Fig.  2: The student MLIPs are Pareto dominant compared to the control models. The student and control MLIPs are trained on the ANI-Al aluminum dataset 23. The accuracy metric, force RMSE, is plotted as a function of the efficiency metrics, namely, time per MD step per atom in (A) and memory per atom in (B). The origin of the plots corresponds to the Pareto optimum solution. The MD simulations are run using the ASE 50 code on a Nvidia A6000 (48GB) GPU.
@@ 2.2 Teacher-Student Training
In the teacher-student training method, a pre-trained teacher model (or models) is used to train a student model to improve speed 47, 43, memory requirements 48, and generalization 46. Knowledge Distillation (KD) 45 is a well known form of teacher-student training. In the original KD publication 45 auxiliary targets are the teacher model’s predicted relative class probabilities (before the application of the softmax operation in the output layer). The auxiliary targets generated by the teacher contain richer information about the structure of the classes in the dataset. The student models were trained on both the ground truth data and the auxiliary targets and performed better than the control models, which have same architecture as the student but are trained only on the ground truth data. Later works have incorporated deeper architecture dependent knowledge 48, 51, 47. KD has been successfully applied to many different architectures such as CNN 52, graph neural networks 47, transformers 48 and more 53. Another variant of the teacher-student training is the “Born Again” (BA) method 46, where the student and teacher models have the same architecture, and the student model can surpass the teacher’s accuracy. Multiple teachers can be used to train a single student model to improve performance 54, 46. Previous works focus on classification tasks on images 45 and text 48, and limit applications to regression tasks on graph structured data 55.
Recently, teacher-student methods have been explored in chemistry for accelerating molecular dynamics 43 and material property prediction 56. In a related work in Ref. 57, the teacher MLIP (trained on QC ground truth) is used to generate synthetic data by running MD under different conditions. The student model is initially pre-trained on the synthetic data then fine-tuned on the QC ground truth. The atomic energy may provide more fine-grained information than the aggregate configuration energy 43, 57.
In the first step of our teacher-student method, we train a complex “teacher” MLIP on a QC dataset, which consists of configuration energy and forces on each atom. This teacher MLIP can generate auxiliary targets, namely per-atom energies. Then, we train the student MLIP on the original QC data and auxiliary targets generated by the teacher MLIP. This economical approach does not require any expensive QC calculations beyond the original dataset needed to train the teacher model, nor exhaustive hyper-parameter tuning.
@@ 2.3 Training Procedure and Loss Function
The $\mathcal{L}_{L_{2}}$ loss term is a regularization of the model weights, which is commonly added to loss functions to reduce over-fitting. The $\mathcal{L}_{\mathrm{R}}$ term, specific to HIPNN, also enhances model stability. The error loss $\mathcal{L}_{\mathrm{err}}$ could be any combination of common metrics, but in our case is an equal weighting between root-mean-squared error (RMSE) and mean-absolute error (MAE) losses:
$\mathcal{L}_{\mathrm{err}}(\hat{V},V)=\mathrm{RMSE}(\hat{V},V)+\mathrm{MAE}(\hat{V},V)$ (3)
where $\hat{V}$ is the model prediction and $V$ is the ground truth target. The weights of each term in the loss are listed in Appendix A.1.
The teacher $\mathcal{T}$ maps the local atomic environments to atomic energies $\epsilon_{i}$ which are summed over to obtain the energy $\hat{E}$. Although $\epsilon_{i}$ is not directly a physical observable, it nonetheless captures important information about how how the local geometry affects the final prediction of the model. Here, we use the teacher MLIP’s atomic energies as the knowledge to be transferred to the student MLIPs.
@@ 2.4 Data
We apply our teacher-student workflow on the ANI-Al dataset, which consists of condensed-phase aluminum geometries, with energies and forces calculated using Density Functional Theory (DFT) 23. The dataset was sampled using an automated active learning framework to generate adequate coverage of the configurational space. The dataset consists of about 6,000 DFT calculations (3 electron PBE ultrasoft pseudo-potential, corresponding to the Al.pbe-n-rrkjus_psl.0.1.UPF potential in Quantum Espresso 58). System volumes were periodic, and contain between $55$ and $250$ aluminum atoms. The thermodynamic conditions span a range of pressures and temperatures. The dataset is available online 59.
@@ 3.1 Pareto Dominant Student MLIPs
Atomistic simulations require accurate and efficient evaluation of energy and forces. Performing MD at large length and time scales is challenging due to the trade-off between accuracy and efficiency of the inter-atomic potential used in the simulation. MLIPs provide a path to perform MD with QC accuracy at dramatically reduced costs relative to methods that use explicit QC solutions to calculate the forces and energies that are used to evolve the simulation forward in time. Here, we show that the teacher-student methods allow us to further improve the accuracy of MLIPs while keeping the computational costs fixed.
Characterizing the accuracy of inter-atomic potentials in absolute terms is challenging 60. Such comparisons should be made against exact numerical solutions or experimental data, which may not be available. The accuracy of MLIPs can be judged against held-out test data from the QC calculations used to develop potential. In this section, we use the force root mean squared error (RMSE) as an accuracy measure. We analyze out-of-sample MD-based accuracy metrics in Sec 3.2.
It is also difficult to define an unambiguous metric of efficiency for MLIPs. An informative measure is the time-per-MD step per atom, which is affected by hardware (CPU/GPU memory and processing speeds) and software (MD library, algorithm used for neighbor list construction, etc.). Therefore, in this article, we use a consistent hardware configuration (a single A6000 GPU with 48 GB of storage) and we use the Atomic Simulation Environment 50 software. We simulate $48,000$ atoms at fixed volume and energy (NVE ensemble) with a time step of 1 fs and a simulation time of 1 ps. Additionally, we can characterize the efficiency of MLIPs based on memory requirements. We record the maximum number of atoms that can be simulated on a single A6000 GPU using the ASE software and use it to calculate the average memory-per-atom.
Figure 2 shows the Pareto plot of accuracy (force RMSE) against the time per MD step per atom and maximum atoms per GPU in panels (A) and (B) respectively. The student models are Pareto dominant with respect to the control MLIPs, i.e., for a given cost (MD speed or memory requirement) the student models have higher accuracy than the control models. Each data point is averaged over four MLIPs initialized with different random seeds.

### v2

# Sakib Matina, Alice E. A. Allena,b,c, Emily Shinkled, Aleksandra Pachalievae, Galen T. Cravena, Benjamin Nebgena, Justin S. Smithf, Richard Messerlya, Ying Wai Lid, Sergei Tretiaka,b,g, Kipton Barrosa,b, and Nicholas Lubbersd

@@
[Uncaptioned image] Teacher-student training improves accuracy and efficiency of machine learning interatomic potentials
Sakib Matina, Alice E. A. Allena,b,c, Emily Shinkled, Aleksandra Pachalievae, Galen T. Cravena, Benjamin Nebgena, Justin S. Smithf, Richard Messerlya, Ying Wai Lid, Sergei Tretiaka,b,g, Kipton Barrosa,b, and Nicholas Lubbersd
[Uncaptioned image] Machine learning interatomic potentials (MLIPs) are revolutionizing the field of molecular dynamics (MD) simulations. Recent MLIPs have tended towards more complex architectures trained on larger datasets. The resulting increase in computational and memory costs may prohibit the application of these MLIPs to perform large-scale MD simulations. Here, we present a teacher-student training framework in which the latent knowledge from the teacher (atomic energies) is used to augment the students’ training. We show that the light-weight student MLIPs have faster MD speeds at a fraction of the memory footprint compared to the teacher models. Remarkably, the student models can even surpass the accuracy of the teachers, even though both are trained on the same quantum chemistry dataset. Our work highlights a practical method for MLIPs to reduce the resources required for large-scale MD simulations.
@@ 1 Introduction
Molecular Dynamics (MD) 1 is ubiquitous in chemistry 2, materials science 3, and drug discovery 4 as well as other fields. Accurate chemical and thermodynamic properties derived from MD rely on accurate interatomic potentials, which parameterize the many-body interactions present between atoms 1, 5. Simulation scales may vary greatly depending on the questions of interest and available resources, and there is a persistent need for greater model efficiency. Traditional classical potentials are very fast, making it possible to perform large-scale simulations of billions of atom 6, or to perform hundreds of millions of integration time-steps for small systems. Recently, there is great demand for interatomic potentials that are more accurate via machine learning models that are trained to reference quantum mechanical force calculations. Here, the goals of efficiency and accuracy can be in conflict. Our paper is concerned with techniques for improving the efficiency of the interatomic potential without sacrificing accuracy.
@@
The gold-standard for computational chemistry is ab initio molecular dynamics, which uses quantum chemistry (QC) methods for accurately calculating interatomic forces from first principles 7, 8. However, the often prohibitive computational cost of QC methods, and its rapid growth with system size, limits the size of the systems that can be simulated. Machine learning interatomic potentials (MLIPs) 9, 10, 11, 12, 13 can be trained on QC datasets to map from an atomic configuration to energy and forces. MLIPs can achieve the chemical accuracy (error $<1\mathrm{kcal/mol}$) of QC simulations 14 at drastically reduced computational costs. Most MLIPs achieve linear scaling with system size by utilizing the approximation that the total predicted energy can be decomposed as a sum of spatially local atom-wise or pair-wise contributions 15, 12, 11, 16. MLIPs have seen explosive growth and have been successfully applied to predicting potential energy surfaces 17, 18, 19, 20, 21, 22, 23, 24, 25, with extensions to a variety of quantities, such as charges 26, 27, 28, 29, 30, and more 31, 32, 33, 34, 35. However, MLIPs only mitigate the cost of QC—they do not eliminate it. Generating training data for MLIPs is costly due to the steep scaling of QC methods in system size $N$, e.g., $\mathcal{O}(N^{7})$ at the CCSD(T) level of theory 7. Therefore any approach which uses data more efficiently can potentially reduce the overall computational costs of MLIPs.
@@
Furthermore, a recent trend in the field is the significant effort to construct foundation model MLIPs 36, 37, 38 with broad chemical coverage39, inspired by the success of large pre-trained models in natural language processing 40. Foundation MLIPs 41, which may be parameterized by up to $10^{9}$ fitting parameters 42, have high computational and memory requirements 43, 44. These increasing computational and memory costs compared to classical force fields 1 can limit the usability of MLIPs for large scale MD simulations 13, 45, 43, 44.
@@
In this manuscript, we introduce a teacher-student training method for MLIP with a central purpose of building the MLIPs with faster inference and lower memory requirements. The teacher-student training is a class of methods 46, 47, 48 to reduce the inference costs of different ML models and more effectively utilize existing datasets. An initial teacher model is trained and then used to augment the training of a student model that may have faster inference 48, 43, 44, smaller memory requirements 49, or better generalization capacity 47. Crucially, this knowledge distillation procedure does not require additional first-principles training data. Instead, auxiliary predictions of the teacher model are used to augment the data used for training of the student model. The innovation in this work is to use the teacher’s local atomic energy predictions as the auxiliary training data. Although these local atomic energies have traditionally been considered a latent feature of the MLIP more 11, the present work highlights that they carry important information. Note that the latent atomic energies provided to the student are far greater in number than the single global QC energy. Thus the student is trained using a significantly larger number of constraints than the teacher. Because the student model will typically have fewer trainable weighs than the teacher model, this approach can yield significant gains for inference speed and memory requirements. We also find that the student model can achieve higher-accuracy than the teacher. An overview of our workflow is represented in Fig 1.
@@
Fig.  1: Teacher-student training for Hierarchically Interacting Particle Neural Network (HIPNN) machine learning interatomic potential. A HIPNN model consists of an input node, a message passing interaction layer and feed-forward regression layers called atom layers. The teacher HIPNN is trained on the quantum mechanical energy and forces data, and generates latent knowledge in the form of atomic energies. This augments the student model training to improve the accuracy. The student HIPNN that has fewer trainable weights, which contributes to faster inference and lower memory requirements. The control HIPNNs have the same number of trainable parameters as the student but are trained only on ground truth data. We show that the student models are more accurate than the control models.
@@ 2.1 Architecture
HIPNN 22 is a message-passing graph neural network 50 that can map atomic configurations to various chemical properties such as energy 22, forces 14, dipoles 28 etc. Similar to many MLIP architectures, the interaction layers of HIPNN allow for mixing of atomic environments between neighbors via message passing to construct the learnable features 22. The local atomic environment of each atom is initially featurized using atomic number (in HIPNN, as a one-hot encoding, although also common is a random embedding) and passed through several layers along with pair-wise displacement vectors between the neighbors with local cut-off to predict the atom energy $\mathcal{\epsilon}_{i}$ for each atom $i$. Automatic differentiation is used to calculate the forces on each atom from the total molecular energy.
@@
Fig.  2: The student MLIPs are Pareto dominant compared to the control models. The student and control MLIPs are trained on the ANI-Al aluminum dataset 23. The accuracy metric, force RMSE, is plotted as a function of the efficiency metrics, namely, time per MD step per atom in (A) and memory per atom in (B). The origin of the plots corresponds to the Pareto optimum solution. The MD simulations are run using the ASE 51 code on a Nvidia A6000 (48GB) GPU. The error bar corresponds to the standard deviation of an ensemble of $4$ models, which differ by random weight initializations.
@@ 2.2 Teacher-Student Training
In the teacher-student training method, a pre-trained teacher model (or models) is used to train a student model to improve speed 48, 43, 44, memory requirements 49, and generalization 47. Knowledge Distillation (KD) 46 is a well known form of teacher-student training. In the original KD publication 46 auxiliary targets are the teacher model’s predicted relative class probabilities (before the application of the softmax operation in the output layer). The auxiliary targets generated by the teacher contain richer information about the structure of the classes in the dataset. The student models were trained on both the ground truth data and the auxiliary targets and performed better than the control models, which have same architecture as the student but are trained only on the ground truth data. Later works have incorporated deeper architecture dependent knowledge 49, 52, 48. KD has been successfully applied to many different architectures such as CNN 53, graph neural networks 48, transformers 49 and more 54. Another variant of the teacher-student training is the “Born Again” (BA) method 47, where the student and teacher models have the same architecture, and the student model can surpass the teacher’s accuracy. Multiple teachers can be used to train a single student model to improve performance 55, 47, 56, 57. Previous works focus on classification tasks on images 46 and text 49, and limit applications to regression tasks on graph structured data 58.
Recently, teacher-student methods have been explored in chemistry for accelerating molecular dynamics 43, 44, physics-constrained data augmentation 59, and material property prediction 60. In a related work in Ref. 61, the teacher MLIP (trained on QC ground truth) is used to generate synthetic data by running MD under different conditions. The student model is initially pre-trained on the synthetic data then fine-tuned on the QC ground truth.
In the first step of our teacher-student method, we train a complex “teacher” MLIP on a QC dataset, which consists of configuration energy and forces on each atom. This teacher MLIP can generate auxiliary targets, namely per-atom energies. The atomic energy may provide more fine-grained information than the aggregate configuration energy 43, 61, 62. Then, we train the student MLIP on the original QC data and auxiliary targets generated by the teacher MLIP. This economical approach does not require any expensive QC calculations beyond the original dataset needed to train the teacher model, nor exhaustive hyper-parameter tuning.
@@ 2.3 Training Procedure and Loss Function
The $\mathcal{L}_{L_{2}}$ loss term is a regularization of the model weights, which is commonly added to loss functions to reduce over-fitting. The $\mathcal{L}_{\mathrm{R}}$ term, specific to HIPNN, also enhances model stability. The error loss $\mathcal{L}_{\mathrm{err}}$ could be any combination of common metrics, but in our case is an equal weighting between root-mean-squared error (RMSE) and mean-absolute error (MAE) losses,
$\mathcal{L}_{\mathrm{err}}(\hat{V},V)=\mathrm{RMSE}(\hat{V},V)+\mathrm{MAE}(\hat{V},V),$ (3)
where $\hat{V}$ is the model prediction and $V$ is the ground truth target. Our loss function uses a sum of the RMSE and MAE errors, similar to the original HIPNN model publications 22, 14. The weights of each term in the loss are listed in Appendix A.1.
The teacher $\mathcal{T}$ maps the local atomic environments to atomic energies $\epsilon_{i}$ which are summed over to obtain the energy $\hat{E}$. Although $\epsilon_{i}$ is not directly a physical observable, it nonetheless captures important information about how the local geometry affects the final prediction of the model 43, 61, 62. Here, we use the teacher MLIP’s atomic energies as the knowledge to be transferred to the student MLIPs.
@@ 2.4 Data
We apply our teacher-student workflow on the ANI-Al dataset 23, which consists of condensed-phase aluminum geometries, with energies and forces calculated using Density Functional Theory (DFT). The authors of reference 23 created the dataset using an automated active learning framework to generate adequate coverage of the configurational space. In this workflow, an ensemble of ANI models 20 were trained to the initial DFT dataset, which consisted of random structures. The main loop of the active learning workflow involves running MD simulations with the ensemble under varying thermodynamic conditions (time-dependent temperatures and density schedules) on boxes of $\approx 50$ to $\approx 250$ atoms. New DFT calculations were performed on configurations where the ensemble disagreement exceeded a predefined threshold. Then the MLIPs were retrained to the expanded dataset. This loop terminates when long MD simulations ($250\mathrm{ps}$) could be performed without identifying any new configurations with high ensemble disagreement. Over 50 generations of models, the final dataset is comprised of about 6,000 DFT calculations with the Perdew-Burke-Ernzerhof functional. The dataset is available online 63. More details are available in the original publication describing the construction of the dataset. 23
@@ 3.1 Pareto Dominant Student MLIPs
Atomistic simulations require accurate and efficient evaluation of energy and forces. Performing MD at large length and time scales is challenging due to the trade-off between accuracy and efficiency of the interatomic potential used in the simulation. MLIPs provide a path to perform MD with QC accuracy at dramatically reduced costs relative to methods that use explicit QC solutions to calculate the forces and energies that are used to evolve the simulation forward in time. Here, we show that the teacher-student method allows us to improve the accuracy of the smaller MLIPs without raising the computational costs at inference time.
Characterizing the accuracy of interatomic potentials in absolute terms is challenging 64. Such comparisons should be made against exact numerical solutions or experimental data, which may not be available. The accuracy of MLIPs can be judged against held-out test data from the QC calculations used to develop potential. In this section, we use the force root mean squared error (RMSE) as an accuracy measure. We analyze out-of-sample MD-based accuracy metrics in Sec 3.2.
It is also difficult to define an unambiguous metric of efficiency for MLIPs. An informative measure is the time-per-MD step per atom, which is affected by hardware (CPU/GPU memory and processing speeds) and software (MD library, algorithm used for neighbor list construction, etc.). Therefore, in this article, we use a consistent hardware configuration (a single A6000 GPU with 48 GB of storage) and we use the Atomic Simulation Environment 51 software. We simulate $48,000$ atoms at fixed volume and energy (NVE ensemble) with a time step of 1 fs and a simulation time of 1 ps. Additionally, we can characterize the efficiency of MLIPs based on memory requirements. We record the maximum number of atoms that can be simulated on a single A6000 GPU using the ASE software and use it to calculate the average memory-per-atom.
Figure 2 shows the Pareto plot of accuracy (force RMSE) against the time per MD step per atom and maximum atoms per GPU in panels (A) and (B) respectively. The student models are Pareto dominant with respect to the control MLIPs, i.e., for a given cost (MD speed or memory requirement) the student models have higher accuracy than the control models. Each data point is averaged over four MLIPs initialized with different random seeds. Relative to the teacher, the student MLIPs can simultaneously achieve more than a factor of 2 speed up in MD simulations at less than half the memory requirements, while sacrificing less than 20% in force accuracy.